The initial speed of a bullet fired from a rifle is 60 m/s.The rifle is fired at the centre of a target 700 m away at the same level as the target.How far above the centre of the target the rifle must be aimed in order to hit the target?

Grade:12

2 Answers

Mallikarjun Maram
9 years ago
The bullet will not hit target for any height about the center of the target, as the maximum range for the given initial speed = u2/g = 3600/10 = 360 m.  Target is at 700 m.  Hence target is out of range for the bullet.
60 m/s is very small speed for bullet.  Please verify the data.
Divya
7 years ago
Yes the question says 630m/s not 60m/s. Can you please provide the solution?
 
__________________________________________________________________________________________________________________________________________________________________________________________________________________